OpenStudy (anonymous):

In some countrys congress, the number of representatives is 30 less than five times the number of senators. there are total of 510 members. find the number of senators and the number of representatives.

OpenStudy (across):

This is a system:\[\text{representatives}=5\cdot\text{senators}-30\]\[\text{representatives}+\text{senators}=510\]

OpenStudy (anonymous):

i did it i keep gettin it wrong thats why im asking for help

OpenStudy (across):

\[\text{representatives}=510-\text{senators}\]\[510-\text{senators}=5\cdot\text{senators}-30\]\[540=6\cdot\text{senators}\] \[\text{senators}=90\]\[\text{representatives}=510-90=420\]

OpenStudy (anonymous):

thanks you so much